OpenStudy (anonymous):

a force of 8.0N acts on a 2.0kg mass for 5.0s a. what is the change in momentum of the mass? b. What is the change in the velocity of the mass?

OpenStudy (kagıtucak):

Lets say its first velocity is 0 m/s . F=ma 8=2.a a=4m/s^2 after 5s its velocity becomes 0+5.4=20 m/s a.momentum=mv at the beginnig m.0=0 at the end mv=2.20=40 40-0=40 kgm/s b. first velocity is 0 final velocity is 20 change is 20-0=20 m/s

OpenStudy (anonymous):

how can we assume its initial is 0?

OpenStudy (kagıtucak):

the question ask the change you can assume any value for initial velocity it doesnt change the answer since it asks the change you can check

OpenStudy (anonymous):

if you don't mind, go through the math with me step by step with 1 and prove that that leads to the exact same answer .

OpenStudy (anonymous):

?